INFRARED RESONANCE ENHANCED PHOTODISSOCIATION SPECTROSCOPY OF METAL ION-BENZENE COMPLEXES

\underline{T.D. JAEGER} and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556.

First row transition metal ion-benzene complexes of the form M^{+}(benzene) _{x} and M^{+}(benzene) _{x} Ar are produced by laser vaporization in a pulsed nozzle source, mass-selected in a time of flight mass spectrometer and excited in the infrared with an OPO/OPA laser system. Multiphoton dissociation occurs by the elimination of whole benzene molecules or argon. The photofragment yield versus wavelength produces IR resonance-enhanced multiphoton photodissociation (IR-REMPD) spectra for these complexes. Vibrational bands observed in the 2700-3300 cm^{−}^{1} region correspond to vibrational modes in the benzene molecules but are shifted from those of free benzene due to the interaction with the metal ion.